CNot secure 1. as follows the Widner Company manufactures twro products Stainless Serving Spoons and Shaunless Serving Forks. The costs and revenues are Saies price 80 4 Total demand for Spoons s 14,000 unts and for Forks is 9000 uvits Machine tme is a scarce resource During the year 541000 machine hous re whle Forks requires 3 machine hours per unit How many units of Spoons and Forks should Widner produce O Opton A O Option B O Osien D o 9.圈@鸟,@-e, et Search Windows ViewSonic
Answer #1

Solution:

Computation of contribution margin per machine hour
Particulars Spoons Forks
Selling Price $150.00 $88.00
Variable cost per unit $80.00 $42.00
Contribution margin per unit $70.00 $46.00
Machine hour per unit 5 3
Contribution margin per machine hour $14.00 $15.33

As contribution margin per machine hour of fork is highest, therefore machine hours should be first utilized in production of Forks.

Total machine hours utilized in production of 9000 forks = 9000*3 = 27000

Nos of spoons to be produced with remaining 27000 machine hours = 27000 / 5 = 5400

Therefore Option D is correct.